cband  instabilities 2011/2012



      PolA of cband has had stability problems 2011-2012 . A dropout in polB of cband was reported on 14jul12.


26oct12:polA drain current jumps around with amp#18

    On 18oct12 cband was warmed up and opened (in the dome). The new amp #18 was replaced with the older amp#19. A bad elbow was found in the dewar.
The plots show the drain current (polA,B) using amp#18 (25sep-18oct12) and amp#19 (pre 25sep and after 18oct12) (.ps)  (.pdf)

processing: x101/121026/

29aug12: replaced cable from bias chassis do dewar

    The cable from the bias chassis to the dewar was replaced on 29aug12 at 8:24 am. Cband is still using the bias chassis from sbh.
A plot was made of the drain currents before and after the cable change (.ps) (.pdf).
processing: x101/120829/

27aug12: cband switched to use sbh bias box

    To debug the polA instabilities, the bias box for cband was replaced with the bias box from sbh on 27aug12.

The plot shows the cband bias drain currents for 27,28aug12 (.ps) (.pdf)

processing: x101/120827/

20aug12: polA total power changes with stage2 drain current rms increase.  (top)

    The output of the cband receiver was monitored for a few hours on 20aug12 to check its stability. The setup was:
The plots show the results of the monitoring (.ps) (.pdf):



31jul12: plot dewar bias values for jul12. (top)

    Jumps in polB were reported on 14jul12 during a vlbi run. The bias monitoring for jul12 was examined to see if it was stable.

The voltage and current monitoring of the dewar  biases from  jul12 is plotted (.pdf):


processing: x101/120731/

31may12: check cband after reinstalled in dome.  (top)

    the cband receiver was brought to the lab on 13may12, worked on , and then reinstalled 30may12. On 31may12 blank sky was tracked with the cband receiver to check the stability. The setups was:

Processing the data:    

The plots show the results of the measurement (.ps) (.pdf):


processing: x101/120531/

09may12:  cross head oscillation causes a gain change. (top)

    Cband was seeing an rf oscillation of about 1% synchronized to the cross head  cycle (1.2hz). High time resolution data was taken with the 25 hz cal running to see if this was an increase of Tsys or a gain change.

 The setup was:

Processing the data:
The plots show the results of the measurements (.ps) (.pdf):


processing: x101/120509/

08-12 may12: Dewar monitoring of Drain voltages and Drain currents. (top)

    A plot was made of the drain current and drain voltage for 08may12-12may12 (.ps) (.pdf):


processing: x101/120510/

cband page